# OOTDiffusion: Outfitting Fusion based Latent Diffusion for Controllable Virtual Try-on

### v1

@@ 4.4 Experimental Results
Table 3: Quantitative results on the Dress Code dataset [33]. The best and second best results are reported in bold and underline, respectively. The * marker refers to the results reported in previous works.
@@
Table 4: Quantitative results of the cross-dataset evaluation. Each model is trained on one of the VITON-HD [6] and Dress Code [33] datasets and evaluated on the other. The best and second best results are reported in bold and underline, respectively. The * marker refers to the results reported in previous works.
@@
Tab. 2 presents the quantitative evaluation results on the VITON-HD dataset [6]. We observe that some GAN-based models like HR-VITON [27] and GP-VTON [52] achieve relatively high SSIM scores, indicating that they are able to retain the structural information of the original images. However, their generated images lack detail fidelity, and thus drop behind ours on LPIPS. The previous LDM-based methods including LaDI-VTON [32] and StableVITON [24] generate more realistic images according to their FID and KID scores, but they fail to restore the detail features due to their lossy feature fusion. In comparison, our OOTDIffusion not only generates realistic outfitted images but also preserves the precise details, and thus substantially outperforms other methods on the other three metrics (LPIPS, FID and KID) while obtaining comparable SSIM scores to the GAN-based methods.
Tab. 3 demonstrates the state-of-the-art performance of our method on the Dress Code dataset [33], which outperforms others on all the metrics for all the garment categories (upper-body/lower-body/dresses), confirming our feasibility in more complicated cases. Note that GP-VTON [52] applies extra data modifications such as background removal and pose normalization to Dress Code, and only provides a part of their test data. Despite this, our OOTDiffusion still achieves the best results on the more challenging original test dataset.
Furthermore, the generalization capability of our method is quantitatively verified by the results of the cross-dataset evaluation listed in Tab. 4. We see that GP-VTON [52] falls far behind other methods on all the metrics since its warping module severely overfits the training data. While our method leads again on all the metrics for the out-of-distribution test dataset. Overall, the observations above (Tabs. 2, 3 and 4) further demonstrate that our OOTDiffusion significantly outperforms other VTON methods in both fidelity and controllability for all kinds of scenarios and conditions.
@@ 4.5 Limitations
Despite the state-of-the-art performance achieved in the image-based virtual try-on task, limitations still exist in our OOTDiffusion which demand further improvement. First, since our model is trained on paired human and garment images, it may fail to get perfect results for cross-category virtual try-on, e.g., to put a T-shirt on a woman in a long dress, or to let a man in pants wear a skirt. This issue can be partially solved in the future by collecting datasets of each model wearing different clothes in the same pose. Another limitation is that some details in the original human image might be altered after virtual try-on, such as muscles, watches and tattoos, etc. This is because the relevant body area is masked and repainted by the diffusion model. Thus more practical pre- and post-processing methods are required for addressing such problems.
@@ 5 Conclusion
In this paper, we present OOTDiffusion, a novel LDM-based network architecture for image-based vitrual try-on. The proposed outfitting UNet efficiently learns the garment features and incorporates them into the denoising UNet via the proposed outfitting fusion process with negligible information loss. Classifier-free guidance for the garment features is enabled by the proposed outfitting dropout in training, which further enhances the controllability of our method. Extensive experiments on high-resolution datasets show our superiority over other VTON methods in both fidelity and controllability, indicating that our OOTDiffusion has broad application prospects for virtual try-on.
@@ Acknowledgements
We sincerely thank our colleagues including Wei Du, Xuping Su, Yilan Ye and Chi Zhang, etc., for kindly supporting and promoting our work.

### v2

@@ 4.4 Experimental Results
Table 3: Quantitative results on the Dress Code dataset [33]. The best and second best results are reported in bold and underline, respectively. The * marker refers to the results reported in previous work.
@@
Table 4: Quantitative results of the cross-dataset evaluation. Each model is trained on one of the VITON-HD [6] and Dress Code [33] datasets, and evaluated on the other. The best and second best results are reported in bold and underline, respectively. The * marker refers to the results reported in previous work.
@@
Tab.˜2 presents the quantitative evaluation results on the VITON-HD dataset [6]. We find that some GAN-based models like HR-VITON [27] and GP-VTON [52] achieve relatively high SSIM scores, indicating that they are able to retain the structural information of the original images. However, their generated images lack detail fidelity, and thus drop behind ours on LPIPS. The previous LDM-based methods including LaDI-VTON [32] and StableVITON [24] generate more realistic images according to their FID and KID scores, but they fail to restore the detail features due to their lossy feature fusion. In comparison, our OOTDiffusion not only generates realistic outfitted images but also preserves the precise details, and thus substantially outperforms other methods on the other three metrics (LPIPS, FID and KID) while obtaining comparable SSIM scores to the GAN-based methods.
Tab.˜3 demonstrates the state-of-the-art performance of our method on the Dress Code dataset [33], which outperforms others on all the metrics for all the garment categories (upper-body/lower-body/dresses), confirming our feasibility in more complicated cases. Note that GP-VTON [52] applies extra data modifications such as background removal and pose normalization to Dress Code, and only provides part of their test data. Despite this, our OOTDiffusion still achieves the best results on the more challenging original test dataset.
Furthermore, the generalization capability of our method is quantitatively verified by the results of the cross-dataset evaluation listed in Tab.˜4. We find that GP-VTON [52] falls far behind other methods on all the metrics since its warping module severely overfits the training data. While our method leads again on all the metrics for the out-of-distribution test dataset. Overall, the observations above (Tabs.˜2, 3 and 4) further demonstrate that our OOTDiffusion significantly outperforms previous VTON methods in both realism and controllability in all kinds of scenarios and conditions.
@@ 4.5 Limitations
Despite the state-of-the-art performance achieved in the image-based virtual try-on task, limitations still exist in our OOTDiffusion which demand further improvement. First, since our models are trained on paired human and garment images, it may fail to get perfect results for cross-category virtual try-on, e.g., to put a T-shirt on a woman in a long dress, or to let a man in pants wear a skirt. This issue can be partially solved in the future by collecting datasets of each person wearing different clothes in the same pose. Another limitation is that some details in the original human image might be altered after virtual try-on, such as muscles, watches or tattoos, etc. The reason is that the relevant body area is masked and repainted by the diffusion model. Thus more practical pre- and post-processing methods are required for addressing such problems.
@@ 5 Conclusion
In this paper, we present OOTDiffusion, a novel LDM-based network architecture for image-based vitrual try-on. The proposed outfitting UNet efficiently learns the garment features and incorporates them into the denoising UNet via the proposed outfitting fusion process with negligible information loss. Classifier-free guidance for the garment features is enabled by the proposed outfitting dropout in training, which further enhances the controllability of our method. Extensive experiments on high-resolution datasets show our superiority over other VTON methods in both realism and controllability, indicating that our OOTDiffusion has broad application prospects for virtual try-on.
@@ Acknowledgements
We sincerely thank our colleagues including Yilan Ye, Bin Fu, Wei Du, Xuping Su, and Chi Zhang, etc., for kindly supporting and promoting our work. Special thanks to Minh-Duc Vo for his helpful advice.